OpenStudy (katherinesmith):

Solve the following system of equations by using the elimination method.

OpenStudy (katherinesmith):

-2x + 9y = -13 6x – 3y = 15

OpenStudy (anonymous):

multiply the equation -2x+9y=-13 with 3 and add to the equation 6x-3y=15, then -6x+27y=-39 + 6x- 3y= 15 => 24y=-24 => y= -1 then x=2

OpenStudy (katherinesmith):

thank you!